OpenStudy (anonymous):

Rectangles A and B are similar rectangles. The length of the diagonal of Rectangle A is 13 inches and the length of the diagonal of Rectangle B is 6.5 inches. What could be the length and width of both Rectangle A and Rectangle B? Show your work. Rectangle A: 5 in x 12 in, Rectangle B: 2.5 in x 6 in Rectangle A: 6.5 in x 14 in, Rectangle B: 3.5 in x 8 in Rectangle A: 4 in x 10 in, Rectangle B: 3 in x 7 in Rectangle A: 7 in x 11 in, Rectangle B: 2 in x 5 in

OpenStudy (anonymous):

tips

OpenStudy (callisto):

ratio of the sides should be the same.

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

rectangle A

OpenStudy (anonymous):

choice 1

OpenStudy (anonymous):

choice 1

OpenStudy (anonymous):

30=30

OpenStudy (callisto):

\[\frac{13}{6.5} = 2\] Check the ratio of the corresponding sides now. For example, for option A, the first pair = 5/2.5 = 2 => true

OpenStudy (anonymous):

so am I right?

OpenStudy (anonymous):

30=30 for choice 1?

OpenStudy (callisto):

How do you get 30? *Note I don't mind if you don't type eh twice. It lengthens the post and possibly makes my computer slow if you continue.

OpenStudy (anonymous):

easy, 5/12=2.5/6

OpenStudy (anonymous):

you see it now?

OpenStudy (anonymous):

12 times 2.5=30 and 5 times 6=30

OpenStudy (anonymous):

did I do something wrong?

OpenStudy (anonymous):

you see what I am saying?

OpenStudy (anonymous):

what do you think, KingGeorge?

OpenStudy (callisto):

Oh... you're equating them and show they are equal. Weird to me :| But option 1 is the right choice. Simply because the corresponding sides share the same ratio. Given:\[\frac{ diagonal \ of \ rectangle \ A}{ diagonal \ of \ rectangle \ B} = \frac{13}{6.5} = 2\]Now, for option 1 \[\frac{Length \ of \ rectangle \ A}{Length \ of \ rectangle \ B} = \frac{5}{2.5} = 2\]\[\frac{Width \ of \ rectangle \ A}{Width \ of \ rectangle \ B} = \frac{12}{6} = 2\] => similar!

OpenStudy (anonymous):

d=so I was right?

OpenStudy (anonymous):

you just cross multiply

OpenStudy (callisto):

I didn't cross multiply. I'm not sure if your method is right, but that doesn't make sense to me..

OpenStudy (anonymous):

is it, KingGeorge?

OpenStudy (kinggeorge):

@luisz your method will not always give the correct solution. In this case, it works because the other options they provided you with are not similar rectangles.

OpenStudy (anonymous):

If you cross multiply the other ratios, they will not equate

OpenStudy (anonymous):

ahh

OpenStudy (kinggeorge):

You can use that method to show similarity between two rectangles, but their ratio of similarity may not be what you wanted.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

y

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

next problem